OpenStudy (anonymous):

quadratic equation 2x^2 - 8x = -7.

OpenStudy (amistre64):

yes it is

OpenStudy (anonymous):

...

OpenStudy (amistre64):

it is a quad eq; what are you spose to do with it?

OpenStudy (anonymous):

i can't figure this one out

OpenStudy (amistre64):

me either ... what is it that we are trying to figure out?

OpenStudy (anonymous):

the answer is going to be -2 \[\pm\] sqrt 2 -2 \[\pm\] 2 sqrt 2 -2 \[\pm\] sqrt 2 ---------------- 2 -2 \[\pm\] sqrt 2/2

OpenStudy (anonymous):

that came out bad

OpenStudy (amistre64):

oh, you want to know the x intercepts

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

using what method? complete square, quad formula, or other?

OpenStudy (anonymous):

quadratic formula ax^2 + bx+ c = 0 then -b +or- sqrt b^2 -4ac ------------------- 2a

OpenStudy (amistre64):

complete the square 2x^2 - 8x = -7 2(x^2 - 4x) = -7 2(x^2 -4x +4 - 4) = -7 2(x^2 -4x +4) - 2(4) = -7 2(x -2)^2 - 8 = -7 2(x -2)^2 = -7 + 8 (x -2)^2 = 1/2 x-2 = +- sqrt(1/2) x =2 +- sqrt(1/2)

OpenStudy (amistre64):

\[ax^2 + bx + c=0\] \[x=\frac{1}{2a}(-b\pm\sqrt{b^2-4ac})\]

OpenStudy (amistre64):

2x^2 -8x +7 = 0 \[x=\frac{1}{2.2}(--8\pm\sqrt{(-8)^2-4.2.7})\] \[x=\frac{1}{4}(8\pm\sqrt{64-56})\] \[x=\frac{1}{4}(8\pm\sqrt{8})\] \[x=\frac{1}{4}(8\pm\sqrt{4.2})\] \[x=\frac{1}{4}(8\pm\sqrt{4}\ \sqrt{2})\] \[x=\frac{1}{4}(8\pm2\sqrt{2})\]

OpenStudy (anonymous):

thanks

OpenStudy (amistre64):

yep